# On the Learning with Augmented Class via Forests

### v1

@@ 4.1 Augmented Gini Impurity for Neural Trees
Generally, a deep neural forest is constructed with a DNN encoder $h\colon\mathcal{X}\rightarrow\mathbb{R}^{d^{\prime}}$ and $m$ differentiable neural trees $\mathcal{D}\mathcal{T}_{1},\mathcal{D}\mathcal{T}_{2},\cdots,\mathcal{D}\mathcal{T}_{m}$. The key point is how to learn differentiable neural trees by incorporating information of augmented class, and our idea is to consider the augmented Gini impurity as an optimization objective for neural trees.
Here, we take neural tree $\mathcal{D}\mathcal{T}_{1}$ as an example, and it is essentially an $l$-layer complete binary tree with hyper-parameter $l\geq 2$. By setting $t=2^{l-1}-1$, we denote by $\{\mathcal{B}_{1},\cdots,\mathcal{B}_{t}\}$ and $\{\mathcal{B}_{t+1},\cdots,\mathcal{B}_{2t+2}\}$ all internal and leaf nodes of $\mathcal{D}\mathcal{T}_{1}$, respectively.
For every internal node $\mathcal{B}_{i}$ ($i\in[t]$), we associate it with a function $f_{i}\colon\mathcal{X}\rightarrow[0,1]$ to represent the probability of being assigned to the left child, and hence the probability of the right child is $1-f_{i}(\bm{x})$. Given encoder $h$, we define
@@
where Sigmoid$(z)=(1+\exp(-z))^{-1}$ and $\bm{w}_{i},b_{i}$ are learned parameters, following Kontschieder et al. (2015).
@@
4:     for Mini-bathces $S^{\prime}_{l}$ and $S^{\prime}_{u}$ from $S_{l}$ and $S_{u}$ do
@@
For every leaf node $\mathcal{B}_{i}$ $(i\in[2t+2]\setminus[t])$, let $\mathcal{I}(\bm{x}\to\mathcal{B}_{i})$ be the event that $\bm{x}$ is assigned to leaf $\mathcal{B}_{i}$, and the probability
@@
where $\mathcal{B}_{i}\swarrow\mathcal{B}_{j}$ and $\mathcal{B}_{j}\searrow\mathcal{B}_{i}$ denote that $\mathcal{B}_{i}$ belongs to the left and right subtrees of $\mathcal{B}_{j}$, respectively. We also denote by
@@
for labeled data $S_{l}$ and unlabeled data $S_{u}$.
@@
For $i\in[2t+2]\setminus[t]$, we define the augmented Gini impurity $\mathcal{G}_{\mathcal{\mathcal{B}}_{i}}(S_{l},S_{u})$ w.r.t. $S_{l}$ and $S_{u}$ as
$\mathcal{G}_{\mathcal{B}_{i}}(S_{l},S_{u})=1-\sum_{k\in[\kappa+1]}\vartheta^{2}_{\mathcal{B}_{i},k}(S_{l},S_{u})\ .$ (8)
Here, $\vartheta_{\mathcal{B}_{i},\kappa+1}(S_{l},S_{u})$ is defined as
$\vartheta_{\mathcal{B}_{i},\kappa+1}(S_{l},S_{u})=\left(1-\sum_{(\bm{x},y)\in S_{l}}\frac{(1-\theta)n_{u}\mu_{\mathcal{B}_{i}}(\bm{x})}{n_{l}n_{\mathcal{B}_{i},u}}\right)_{+}\ ;$
and for $k\in[\kappa]$, we define $\vartheta_{\mathcal{B}_{i},k}(S_{l},S_{u})$ as follows:
@@
This definition essentially follows the augmented Gini impurity of decision trees in Definition 1, whereas deep neural trees consider probabilistic assignment of instances to nodes while Definition 1 focuses on deterministic assignment. Here, $\vartheta_{\mathcal{B}_{i},k}(S_{l},S_{u})$ estimates the probability of the $k$-th class under the condition $\mathcal{I}(\bm{x}\to\mathcal{B}_{i})$. Based on Assumption 1, we have
@@
Similar to Theorem 1, we also study the optimal squared loss $\mathcal{L}^{*}_{\mathcal{B}_{i}}$ under the condition $\mathcal{I}(\bm{x}\to\mathcal{B}_{i})$ as follows:
@@
where $\tilde{\bm{y}}\in\mathbb{R}^{\kappa+1}$ is the one-hot encoding of $y\in[\kappa+1]$. Based on Lemma 2 and Assumption 1, we have
@@ 4.2 Deep Neural LACForest
In the training phase, we begin with random initializations on encoder $h$ and neural trees $\mathcal{D}\mathcal{T}_{1},\cdots,\mathcal{D}\mathcal{T}_{m}$, and update the encoder and neural trees iteratively. In each iteration, we receive two mini-batches of labeled data $S^{\prime}_{l}$ and unlabeled data $S^{\prime}_{u}$ sampled randomly from $S_{l}$ and $S_{u}$, respectively. From Definition 2, we introduce
@@
$\omega_{i,j}(S^{\prime}_{u})=\sum_{\bm{x}\in S^{\prime}_{u}}{\mu_{\mathcal{B}_{i,j}}(\bm{x})}/{|S^{\prime}_{u}|}\ .$
@@
Notice that the optimization objective $\mathcal{L}$ in Eqn. (11) is differentiable w.r.t. the parameters of feature encoder $h$ and neural trees $\mathcal{D}\mathcal{T}_{1},\cdots,\mathcal{D}\mathcal{T}_{m}$, and we could update the model with stochastic gradient descent directly.
@@
where $\mathcal{D}\mathcal{T}_{i,k}(\bm{x})$ is the probability of the $k$-th class w.r.t. neural tree $\mathcal{D}\mathcal{T}_{i}$, and could be written as
@@
Algorithm 2 presents a detailed description of our deep neural LACForest approach. In practice, we should adopt relatively large batch sizes for mini-batches $S^{\prime}_{l}$ and $S^{\prime}_{u}$, since $\mathcal{G}_{\mathcal{B}_{i,j}}(S^{\prime}_{l},S^{\prime}_{u})$ in Eqn. (10) converges to the optimal squared loss in the rate of $O(1/\sqrt{\min\{|S^{\prime}_{l}|,|S^{\prime}_{u}|\}})$ from Theorem 2.
Related works
Zhou and Chen Zhou and Chen (2002) proposed the problem of class-incremental learning via a few labeled augmented instances. Fink et al. Fink et al. (2006) learned multiple binary classifiers for this problem and Topalis and Polikar Topalis and Polikar (2008) considered voting classifiers. Recent years have witnessed increasing attention on the design of practical algorithms for this problem Li and Hoiem (2017); Rebuffi et al. (2017); Yan et al. (2021); Zou et al. (2022); Zhou et al. (2024). These methods can not be applied to our learning scenario directly, since we have no access to any labeled augmented instances.
Another relevant problem is open-set recognition in computer vision. Scheirer et al. Scheirer et al. (2013) introduced open space risk to penalize predictions outside the support of training data. Along this line, various approaches have been developed based on open space risk Scheirer et al. (2014), extreme value theory Bendale and Boult (2016); Rudd et al. (2017), nearest neighbors Mendes Júnior et al. (2017) and generative neural networks Ge et al. (2017); Neal et al. (2018); Chen et al. (2021). Most studies are based on strong geometric assumptions, while our work relaxes such assumptions.
Figure 2: Experimental comparisons of AUC on the detection of augmented class for compared methods.
@@ 5.1 Evaluation of our LACForest approach
We compare with the state-of-the-art approaches as follows:
  • GLAC: Learning with augmented class via unbiased risk estimation and linear models Shu et al. (2023);
  • EULAC: Learning with augmented class via unbiased risk estimation and kernel methods Zhang et al. (2020);
  • PAC-iForest: Open category detection based on iForest Liu et al. (2018);
  • EVM: Open set recognition based on extreme value theory and radial basis functions Rudd et al. (2017);
  • OSNN: Open set recognition based on nearest neighbors Mendes Júnior et al. (2017);
  • LACU-SVM: Learning with augmented class based on low-density assumption and SVM Da et al. (2014);
  • OVR-SVM: One-vs-rest method for multi-class learning based on SVM Rifkin and Klautau (2004);
For each dataset, we randomly select half of classes as the augmented class and the rest as known classes, following Zhang et al. (2020). We then randomly sample 500 examples of known classes as labeled data $S_{l}$, and 1000 instances as unlabeled data $S_{u}$ and 100 instances as testing data. We take $\theta=0.5$ in Eqn. (1), and more experimental settings could be found in Appendix B.1. The performance is evaluated by 10 trials of random selections of augmented class, and with 10 times of random data sampling. The average test accuracies are obtained over these 100 runs, as shown in Table 2.
@@ 5.2 Evaluation of our deep neural LACForest
We take a three-layer convolutional neural network as the backbone neural network on datasets mnist, fmnist and kuzushiji, and consider VGG16 Simonyan and Zisserman (2015) on svhn and cifar10, as done in Zhang et al. (2020); Shu et al. (2023). We randomly select four classes as augmented class and take the rest as the known classes, and set $\theta=0.4$ similarly to Shu et al. (2023).
We take average test accuracies over 5 random selections of augmented class as our performance measure, and the experimental comparisons are shown in Table 3. It is clear that our deep neural takes better performance than ARPL, G-Openmax, OSRCI, Openmax and Softmax-T, due to some additional geometric assumptions over those methods. For Deep-GLAC and Deep-EULAC, our method achieves better and comparable performance in most times, and an intuitive explanation is that our approach could effectively explore augmented class from each local region during tree partitions.
Parameter Influence
We analyze the influence of various parameters on several datasets, and the trends are similar for other datasets. Figure 3 shows that our LACForest generally gets better performance under different proportions $\theta\in[0.3,0.8]$ for augmented data. Figure 4 shows the performance of LACForest w.r.t. different sizes of unlabeled data, and our LACForest takes better and stable performance as the increase of unlabeled data, which is consistent with our Theorem 1. For deep neural LACForest, Figure 5 shows that our approach is not sensitive to trade-off parameter $\lambda_{\text{ce}}$ and generally works well for $\lambda_{\text{ce}}\in[0.2,2]$. Figure 6 shows the influence of depth of deep neural trees, and our method takes stable results when the tree depth $l\geq 5$.
@@ 6 Conclusion
This work studies learning with augmented class via forests, where an augmented class may appear in testing data yet not in training data. We introduce the augmented Gini impurity as a new splitting criterion by incorporating information of augmented class into trees’ splitting. We then develop the approach on Learning with Augment Class via Forests, which constructs shallow forests based on augmented Gini impurity and further splits forests with pseudo-labeled augmented instances for better performance. We also explore deep neural forests with a new optimization objective via our augmented Gini impurity. We validate the effectiveness of our methods both empirically and theoretically. An interesting future work is to exploit our methods under other learning settings such as streaming datasets and multiple augmented classes.

### v2

@@ 4.1 Augmented Gini Impurity for Neural Trees
A deep neural forest is constructed with a DNN encoder $h$ and $m$ differentiable neural trees $\mathcal{D}\mathcal{T}_{1},\mathcal{D}\mathcal{T}_{2},\cdots,\mathcal{D}\mathcal{T}_{m}$. The key point is how to learn differentiable neural trees by incorporating information of augmented class, and our idea is to consider augmented Gini impurity as an optimization objective for neural trees.
We take neural tree $\mathcal{D}\mathcal{T}_{1}$ as an example, and it is essentially an $l$-layer complete binary tree with $l\geq 2$. We denote by $\{\mathcal{B}_{1},\cdots,\mathcal{B}_{t}\}$ and $\{\mathcal{B}_{t+1},\cdots,\mathcal{B}_{2t+2}\}$ all internal and leaf nodes of $\mathcal{D}\mathcal{T}_{1}$, respectively, where $t=2^{l-1}-1$.
For every internal node $\mathcal{B}_{i}$ ($i\in[t]$), we associate with a function $f_{i}\colon\mathcal{X}\rightarrow[0,1]$ to represent the probability of being assigned to the left child, and hence the probability of the right child is $1-f_{i}(\bm{x})$. Given encoder $h$, we define
@@
where sigmoid$(z)=(1+\exp(-z))^{-1}$ and $\bm{w}_{i},b_{i}$ are learned parameters, as in Kontschieder et al. (2015).
@@
4:     for mini-batches $S^{\prime}_{l}$ and $S^{\prime}_{u}$ from $S_{l}$ and $S_{u}$ do
@@
For every leaf node $\mathcal{B}_{i}$ $(i\in[2t+2]\setminus[t])$, let $\mathcal{I}(\bm{x}\to\mathcal{B}_{i})$ be the event that $\bm{x}$ is assigned to leaf $\mathcal{B}_{i}$, and denote by
@@
where $\mathcal{B}_{i}\swarrow\mathcal{B}_{j}$ and $\mathcal{B}_{j}\searrow\mathcal{B}_{i}$ represent that $\mathcal{B}_{i}$ belongs to the left and right subtrees of $\mathcal{B}_{j}$, respectively. We also denote by
@@
for unlabeled data $S_{u}$ and labeled data $S_{l}$, respectively.
@@
For $i\in[2t+2]\setminus[t]$, we define the augmented Gini impurity w.r.t. $S_{l}$ and $S_{u}$ as
$\mathcal{G}_{\mathcal{B}_{i}}(S_{l},S_{u})=1-\sum_{k\in[\kappa+1]}\vartheta^{2}_{\mathcal{B}_{i},k}(S_{l},S_{u})\ ,$ (8)
where $\vartheta_{\mathcal{B}_{i},\kappa+1}(S_{l},S_{u})$ is defined as
$\vartheta_{\mathcal{B}_{i},\kappa+1}(S_{l},S_{u})=\left(1-\frac{(1-\theta)n_{u}n_{\mathcal{B}_{i},l}}{n_{l}n_{\mathcal{B}_{i},u}}\right)_{+}\ ,$
and for $k\in[\kappa]$, $\vartheta_{\mathcal{B}_{i},k}(S_{l},S_{u})$ is given by
@@
Table 1: Details of datasets.
This definition follows the augmented Gini impurity of decision trees in Definition 1, whereas deep neural trees consider probabilistic assignment of instances to nodes while the latter focuses on deterministic assignment. Essentially, $\vartheta_{\mathcal{B}_{i},k}(S_{l},S_{u})$ estimates the probability of the $k$-th class under the condition $\mathcal{I}(\bm{x}\to\mathcal{B}_{i})$. Based on Assumption 1, we have
@@
Denote by $\mathcal{L}^{*}_{\mathcal{B}_{i}}$ the optimal squared loss under $\mathcal{I}(\bm{x}\to\mathcal{B}_{i})$:
@@
where $\tilde{\bm{y}}\in\mathbb{R}^{\kappa+1}$ is the one-hot encoding of $y\in[\kappa+1]$. Based on Assumption 1 and Lemma 2, we have
@@ 4.2 Deep Neural LACForest
In the training phase, we randomly initialize encoder $h$ and neural trees $\mathcal{D}\mathcal{T}_{1},\cdots,\mathcal{D}\mathcal{T}_{m}$, and update the encoder and neural trees iteratively. In each iteration, we receive two mini-batches of labeled data $S^{\prime}_{l}$ and unlabeled data $S^{\prime}_{u}$ sampled randomly from $S_{l}$ and $S_{u}$, respectively. We introduce
@@
$\omega_{i,j}(S^{\prime}_{u})=\frac{1}{|S^{\prime}_{u}|}\sum_{\bm{x}\in S^{\prime}_{u}}{\mu_{\mathcal{B}_{i,j}}(\bm{x})}\ .$
@@
Notice that the optimization objective $\mathcal{L}$ in Eqn. (11) is differentiable w.r.t. the parameters of feature encoder $h$ and neural trees $\mathcal{D}\mathcal{T}_{1},\cdots,\mathcal{D}\mathcal{T}_{m}$, and we could update the entire model with stochastic gradient descent directly.
@@
where $\mathcal{D}\mathcal{T}_{i,k}(\bm{x})$ shows the probability of the $k$-th class w.r.t. neural tree $\mathcal{D}\mathcal{T}_{i}$, i.e.,
@@
Algorithm 2 presents the detailed description of our deep neural LACForest approach. In experiments, we should take relatively large batch sizes for mini-batches $S^{\prime}_{l}$ and $S^{\prime}_{u}$, since $\mathcal{G}_{\mathcal{B}_{i,j}}(S^{\prime}_{l},S^{\prime}_{u})$ in Eqn. (10) converges to the optimal squared loss in the rate of $O(1/\sqrt{\min\{|S^{\prime}_{l}|,|S^{\prime}_{u}|\}})$ from Theorem 2.
4.3 Related Works
Zhou and Chen Zhou and Chen (2002) introduced the problem of class-incremental learning via a few labeled augmented instances. Fink et al. Fink et al. (2006) learned multiple binary classifiers for this problem and Topalis and Polikar Topalis and Polikar (2008) considered voting classifiers. Recent years have witnessed increasing attention on the design of practical algorithms for this problem Li and Hoiem (2017); Rebuffi et al. (2017); Yan et al. (2021); Zou et al. (2022); Zhou et al. (2024). These methods can not be applied to our learning scenario directly, since we have no access to any labeled augmented instances.
Another relevant problem is open-set recognition in computer vision. Scheirer et al. Scheirer et al. (2013) introduced open space risk to penalize predictions outside the support of training data. Along this line, various approaches have been developed based on open space risk Scheirer et al. (2014), extreme value theory Bendale and Boult (2016); Rudd et al. (2017), nearest neighbors Mendes Júnior et al. (2017) and generative neural networks Ge et al. (2017); Neal et al. (2018); Chen et al. (2021). Those studies are strongly based on some geometric assumptions.
Figure 2: Experimental comparisons of AUC on the detection of augmented class.
@@ 5.1 Evaluation of Our LACForest Approach
For our LACForest approach, we compare with seven state-of-the-art approaches: GLAC Shu et al. (2023), EULAC Zhang et al. (2020), PAC-iForest Liu et al. (2018), EVM Rudd et al. (2017), OSNN Mendes Júnior et al. (2017), LACU-SVM Da et al. (2014) and OVR-SVM Rifkin and Klautau (2004). More details on these approaches could be found in Appendix B.1.
For each dataset, we randomly select half of classes as the augmented class with the rest as known classes, following Zhang et al. (2020). We then randomly sample 500 examples of known classes as labeled data $S_{l}$, and 1000 instances as unlabeled data $S_{u}$ and 100 instances as testing data. We take $\theta=0.5$ in Eqn. (1), and more experimental settings could be found in Appendix B.1. The performance is evaluated by 10 trials of random selections of augmented class, and with 10 times of random data sampling. The average test accuracies are obtained over these 100 runs, as shown in Table 2.
@@ 5.2 Evaluation of Our Deep Neural LACForest
We take a three-layer convolutional neural network as the backbone neural network on mnist, fmnist and kuzushiji, and consider VGG16 Simonyan and Zisserman (2015) on svhn and cifar10, as done in Shu et al. (2023). We randomly select four classes as the augmented class and take the rest as known classes, and set $\theta=0.4$ similarly to Shu et al. (2023).
We take average test accuracies over 5 random selections of augmented class as our performance measure, and the experimental comparisons are shown in Table 3. It is clear that our approach takes better performance than ARPL, G-Openmax, OSRCI, Openmax and Softmax-T, due to some additional geometric assumptions over those methods. For Deep-GLAC and Deep-EULAC, our method achieves better and comparable performance in most times, and an intuitive explanation is that our approach could effectively explore augmented class from each local region during tree partitions.
Parameter Influence. We analyze the influence of various parameters on several datasets, and the trends are similar for other datasets. Figure 3 shows that our LACForest gets better performance under different proportions $\theta\in[0.3,0.8]$ for augmented data. Figure 4 shows the performance with different sizes of unlabeled data, where our LACForest achieves better and stable performance with the increase of unlabeled data, in consistency with Theorem 1. For deep neural LACForest, Figure 5 shows that our approach is insensitive to parameter $\lambda_{\text{ce}}$ and generally works well for $\lambda_{\text{ce}}\in[0.2,2]$. Figure 6 shows the influence of the depth of neural trees, and our method takes stable results when the tree depth $l\geq 5$.
@@ 6 Conclusion
This work studies learning with augmented class via forests, where an augmented class may appear in testing data yet not in training data. We introduce the augmented Gini impurity as a new splitting criterion by incorporating information of augmented class during tree construction. We develop the approach on Learning with Augmented Class via Forests, which constructs shallow forests based on augmented Gini impurity and further splits forests with pseudo-labeled augmented instances. We also explore deep neural forests with a new optimization objective via our augmented Gini impurity. We validate the effectiveness of our methods both empirically and theoretically. An interesting future direction is to exploit our methods under other learning settings such as streaming datasets and multiple augmented classes.